OpenStudy (anonymous):

sin(sec^-1(x+3/4)) help, I don't even know where to start with this one.

OpenStudy (anonymous):

ugliness for sure draw a triangle first

OpenStudy (anonymous):



OpenStudy (anonymous):

ok did the triangle... but bec its sec, wasnt sure how to label it

OpenStudy (anonymous):

find the third side via pythagoras

OpenStudy (anonymous):

thinking real basic as secant equals hypotenuse over adjacent" that is how i labelled it

OpenStudy (anonymous):

i'm sure you meant (x+3)/4 yes?

OpenStudy (anonymous):

but before we go further, is it \[x+\frac{3}{4}\] or is it \[\frac{x+3}{4}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

what @xxferrocixx said if it is the second then label differently

OpenStudy (anonymous):

put the hypotenuse as \(x+3\) and the adjacent as \(4\)

OpenStudy (anonymous):

draw tool is no longer working for me, but i assume you get the idea

OpenStudy (anonymous):

did you ever get help with that last problem btw? the one where you had to rotate about x = -1?

OpenStudy (anonymous):

ok so for the other side I got: \[\sqrt{x ^{2}+6x+7}\]

OpenStudy (anonymous):

I did I struggled with it... but I guess with time I will become more confident with this stuff

OpenStudy (anonymous):

if that is \[\sqrt{(x+3)^2-4^2}\]then yes

OpenStudy (anonymous):

hmm but i think it is not

OpenStudy (anonymous):

oops let me do it over

OpenStudy (anonymous):

should be \[\sqrt{x^2+6x-7}\] i think you subtracted wrong and that is not the final answer in any case, you have to put that over the hypotenuse

OpenStudy (anonymous):

thats the answer? that simple?

OpenStudy (anonymous):

wow i guess it is that simple

OpenStudy (zarkon):

It is a little more complicated than that (just a little)

OpenStudy (anonymous):

that is not the answer, that is the "opposite" side the answer for \[\sin\left(\sec^{-1}(\frac{x+3}{4})\right)\]would be \[\frac{\sqrt{x^2+6x-7}}{x+3}\] as in "opposite over hypotenuse"

OpenStudy (anonymous):

the most complicated part is writing it!

OpenStudy (zarkon):

what do you get if you plug in -10 for x

OpenStudy (anonymous):

i give up what?

OpenStudy (zarkon):

not the answer if you use the function

OpenStudy (zarkon):

you need \[\frac{\sqrt{x^2+6x-7}}{|x+3|}\]

OpenStudy (zarkon):

the range of the arcsec is between 0 and pi and the sin on 0 to pi is positive

OpenStudy (anonymous):

ooh jeez it is always something

OpenStudy (anonymous):

OK ... well thank you satellite, for your help, according to the software, the answer you helped me with is right

OpenStudy (anonymous):

if you want, repost the last problem we can do it

OpenStudy (anonymous):

@Zarkon yeah i see my mistake too if i did it the real way the denominator would be \[\sqrt{(x+3)^2}\]

OpenStudy (anonymous):

@itsmecaro tell the software people you found a mistake (even though it was @Zarkon that found it) maybe they will offer you a reward

myininaya (myininaya):

Like one of those Christmas tree Debbie cakes?

OpenStudy (anonymous):

hahaha a reward??? right they will just blow me off

OpenStudy (anonymous):

i was thinking more along the lines of unused halloween kit kats

OpenStudy (anonymous):

ewwww

myininaya (myininaya):

What? They are unused?

OpenStudy (anonymous):

saw someone returning a whole bag to target earlier today they said they have to throw them out

OpenStudy (anonymous):

i dont know. ive heard some creepy things where they inject stuff in the chocolates

OpenStudy (anonymous):

it's called fat`

OpenStudy (perl):

You can solve this using trig identities. sin (arcsec((x+3)/4)) Let y = arcsec (x+3)/4 sec y = (x+3)/4 cosy = 4/(x+3) Trig identity: sin^2(y) + cos^2(y) = 1 sin^2(y) = 1 - cos^2(y) sin (y) = + - sqrt( 1 - cos^2(y)) sin(y) = + - sqrt[1 - (4/(x+3))^2]

OpenStudy (zarkon):

you don't get a \(\pm\) because \(\sin(y)\ge0\) (because \(0\le y\le \pi\))

OpenStudy (perl):

was that given in the problem , or you derived that

OpenStudy (perl):

i guess because the range of arcsec is positive

OpenStudy (zarkon):

the range of the arcsec(x) is \([0,\pi/2)\cup(\pi/2,\pi]\)

OpenStudy (perl):

ok and sine of that is positive.

OpenStudy (zarkon):

yes

OpenStudy (perl):

that makes sense. how did you know that satellite needed an absolute value. its not obvious with the 'triangle' method

OpenStudy (perl):

if you draw a triangle, i mean

OpenStudy (zarkon):

because of the domains and ranges of arcsec(x) and sin(x)

OpenStudy (zarkon):

you can also get it from the triangle...which satellite73 mentions

OpenStudy (perl):

it seems to work algebraically though. also the hypotenuse is x+3, so that should always be positive, so i guess | x + 3 | makes sense. sin (arcsec((x+3)/4)) Let y = arcsec (x+3)/4 sec y = (x+3)/4 cosy = 4/(x+3) Trig identity: sin^2(y) + cos^2(y) = 1 sin^2(y) = 1 - cos^2(y) sin (y) = + sqrt( 1 - cos^2(y)) sin(y) = + sqrt[1 - (4/(x+3))^2] sin(y) = sqrt[ (x+3)^2 - 4^2 ) / (x+3)^2] sin(y) = sqrt( x^2 + 6x -7 ) / sqrt((x+3)^2) sin(y) = sqrt(x^2 + 6x -7) / |x+3|

OpenStudy (perl):

satellite says "yeah i see my mistake too if i did it the real way the denominator would be sqrt((x+3)^2) " I can only imagine satellite meant by the 'real' way, using the trig identities.

OpenStudy (anonymous):

since you guys are math genius' can u help me out with this next one?